The sum of two numbers is 90. The larger number is 9 less than 3 times the smaller number. Find the smaller number.
Answer by wgomero(28)

small # = x
large # = 3x - 9
equation = x + 3x - 9 = 90
4x = 99
x=99/4
x=24.75
y = 3(24.75) - 9 = 65.25
24.75 + 65.25 = 90